OpenStudy (cas_fangirl_14):

Will give a medal. Solve for x: 2(x+7)-3(2x-4)=-18 A. x=5 B. x=11 C. x=-11 D. x=-5

Vocaloid (vocaloid):

distribute first 2(x+7) = ?

OpenStudy (cas_fangirl_14):

2x+2*7 2x+14?

Vocaloid (vocaloid):

good now, what is 3(2x-4)?

OpenStudy (cas_fangirl_14):

3*2x+3(-4) 3*2x-3*4 6x-12?

Vocaloid (vocaloid):

good, now on the left side we have 2x + 14 - (6x-12) which equals 2x + 14 - 6x + 12 we simplify that to get -4x + 26 = -18 know where to go from there?

OpenStudy (cas_fangirl_14):

Do we divide -4 from each side of the equation?

Vocaloid (vocaloid):

we could, if you wanted to. that would give us x - 6.5 = 4.5 know where to go now?

OpenStudy (cas_fangirl_14):

x-6.5=4.5 x=6.5+4.5 x=11?

Vocaloid (vocaloid):

great! x = 11 is your answer

OpenStudy (cas_fangirl_14):

Thanks :)